# Three-dimensional gravity-capillary standing waves: computation, resonance and instability

### v1

@@ 1 Introduction
Two-dimensional standing water waves have been extensively investigated through rigorous analysis, weakly nonlinear theory, fully nonlinear simulations, and laboratory experiments. Plotnikov & Toland [plotnikov2001nash] and Iooss et al. [iooss2005standing] proved the existence of small-amplitude gravity standing waves in finite and infinite depth, respectively. The Stokes-type expansions of gravity standing waves were derived by Rayleigh [Rayleigh1915deep] and Penney & Price [penney1952some] for deep water, by Tadjbaksh & Keller [tadjbakhsh1960standing] for finite-depth water, and later extended to include surface tension by Concus [concus1962standing] and Vanden-Broeck [vanden1984nonlinear]. In recent years, direct numerical computations on two-dimensional standing waves have received more attention in studying the branching phenomenon [okamura1985instability, mercer1994form, smith1999branching, wilkening2012overdetermined, shelton2023structure], the mechanisms of instability [mercer1992standing, bryant1994different, wilkening2019harmonic], and the long-standing conjecture on the limiting configuration of gravity standing waves [mercer1992standing, schultz1998highly, okamura1998enclosed, okamura2003standing, okamura2010almost, wilkening2011breakdown, wilkening2012overdetermined]. Three principal numerical approaches have been established: (i) spatio-temporal collocation methods based on the Fourier and other spectral expansions [vanden1981numerical, tsai1987numerical, tsai1994numerical, bryant1994different, bryant1995water, okamura2003standing], (ii) semi-analytic series-expansion methods [schwartz1981semi, rottman1982steep, abassi2025semi], and (iii) time-dependent methods (usually via the boundary-integral formulation) combined with Newton or nonlinear least-squares solvers [mercer1992standing, schultz1998highly, wilkening2011breakdown, wilkening2012overdetermined, shelton2023structure]. The last approach is most commonly used owing to its relatively low computational cost, although the time integration may suffer from numerical instabilities, aliasing errors, and filtering errors. In contrast, the spatio-temporal collocation methods avoid these issues and guarantee exact temporal periodicity owing to the Fourier expansion in time. When the system is numerically stiff due to surface tension, this approach is particularly attractive, despite the heavy computational cost in solving large-scale algebraic systems. Such an issue had limited the application of spatio-temporal collocation methods in early years, but is largely alleviated over the past decades with a rapid growth of computational power.
@@ 2 Formulations of deep-water gravity-capillary waves
where $\mathcal{N}_{i}$ represents the DNO-involved $i$-th order nonlinear term whose expression can be found in Appendix 1. Wang & Milewski [wang2012dynamics] and Wang et al. [wang2014computation] employed a cubic ($n=3$) and a quintic ($n=5$) truncations (denoted by cubic model and quintic model hereafter) to study gravity-capillary and flexural-gravity solitary waves, respectively. They showed that these models are quite accurate compared with the original water-wave formulation, even when solutions have large amplitudes. In the following sections, we investigate three-dimensional gravity-capillary standing waves in rectangular or square basins (see figure 1) by searching for triply periodic (in the $x,y$ and $t$ directions) solutions of the cubic and quintic models.
Figure 1: A schematic of gravity-capillary standing waves in a square basin.
@@ 3.2 Spatio-temporal symmetries
    $\eta(x,y,T/4-t^{\prime})$ $=\eta(x\pm L_{1}/2,y,T/4+t^{\prime})=\eta(x,y\pm L_{2}/2,T/4+t^{\prime}),$ (22)
    $\varphi(x,y,T/4-t^{\prime})$ $=-\varphi(x\pm L_{1}/2,y,T/4+t^{\prime})=-\varphi(x,y\pm L_{2}/2,T/4+t^{\prime}).$ (23)
@@
    $\eta(x,y,T/4-t^{\prime})$ $=\eta(x\pm L_{1}/2,y\pm L_{2}/2,T/4+t^{\prime}),$ (24)
    $\varphi(x,y,T/4-t^{\prime})$ $=-\varphi(x\pm L_{1}/2,y\pm L_{2}/2,T/4+t^{\prime}).$ (25)
@@
    $\eta(-x,y,T/4)$ $=\eta(x\pm L_{1}/2,y,T/4)=\eta(x,y\pm L_{2}/2,T/4)=\eta(x,-y,T/4),$ (36)
    $\varphi(-x,y,T/4)$ $=-\varphi(x\pm L_{1}/2,y,T/4)=-\varphi(x,y\pm L_{2}/2,T/4)=\varphi(x,-y,T/4).$ (37)
@@ 3.3 Numerical implementation
$x_{i}$ $=\frac{2iL_{1}}{I+1}-L_{1},\,\quad i=1,2,\cdots,I+1,$ (40)
$y_{j}$ $=\frac{2jL_{2}}{J+1}-L_{2},\quad j=1,2,\cdots,J+1,$ (41)
$t_{n}$ $=\frac{nT}{N+1},\quad\qquad n=1,2,\cdots,N+1,$ (42)
@@
All together, these give rise to (I/4+1)(J/2+1)N/2J/4=O(IJN/16)(I/4+1)(J/2+1)N/2-J/4=O(IJN/16) unknowns for $\eta$ and $\varphi$. For Case II and Case III, it is convenient to let $I=J$. The total number of unknowns for $\eta$ and $\varphi$ are reduced to $(I/4+1)^{2}N/2=O(I^{2}N/32)$ for Case II and $(I/4+1)(I/2+1)N/2=O(I^{2}N/16)$ for Case III. A schematic is drawn in figure 6 to show the reduced computational domain for all three cases. (10)-(11) give rise to the same number of equations after using the spatio-temporal symmetries. To close the system, we use the following wave-amplitude equation
@@ 4.2 Standing waves and bifurcations
We first show representative Case II standing waves with $k=l=1$. In contrast to the gravity standing waves featuring pyramidal-shaped crests, gravity-capillary standing waves exhibit rounded crests and troughs. Figure 9 displays a solution with $H=0.7,\omega=1.867$, calculated using $128\times 128\times 128$ grid points. Panels (a)-(c) correspond to the wave profiles at $t=0,T/4$ and $T/2$, respectively. At $t=T/4$, the surface develops multiple crests and troughs with the maximum wave amplitude decreasing to $0.16$. Figure 11 exhibits the energy distribution among the kinetic, gravitational, and surface-tension parts over one temporal period. The kinetic energy vanishes every half period, while the potential energy reaches its minimum, $0.497$, at $t=T/4,3T/4,\cdots$. The total energy remains at a constant level of $7.661$ perfectly, with the maximum relative error being $7\times 10^{-8}$.
@@
Figure 12: (a)-(e) Surface profiles of a resonant Case II standing wave with $k=l=1,H=0.33$ at $t=0,T/16,3T/16,7T/32$ and $T/4$. (f) Amplitudes of the first $100$ Fourier coefficients for $\eta(x,y,0)$.
A common feature for two-dimensional standing waves is the existence of various nonlinear resonant solutions [smith1999branching, wilkening2011breakdown, wilkening2012overdetermined]. We observe a similar phenomenon in three dimensions. Figure 12 shows a representative Case II resonant standing wave with $k=l=1,H=0.33,\omega=2.0192$, calculated using $128\times 128\times 128$ grid points. In contrast to the smooth surface profiles shown in figure 9, the resonant standing wave exhibits complex surface wrinkles. A local hump and dip alternately emerge on top of the crest at a faster oscillating frequency relative to the fundamental Fourier mode, as shown in panels (a)-(c). The surface develops a volcano-shaped appearance with four saddles along the diagonals at $t=7T/32$, as shown in panel (d). At $t=T/4$, the wave amplitude drops to $0.0247$ approximately, and the surface profile shows a similar pattern to that in figure 9 (b) but has much flatter crests and four sharper spikes at $(\pm\pi/2,\pm\pi/2)$. Panel (f) shows the Fourier amplitudes associated with $\cos(nx)\cos(my)$ mode for $\eta(x,y,0)$. Note that all Fourier modes with odd-valued sum $n+m$ vanish, in consequence of (22) and (23). Owing to the eightfold symmetry, the spectrum is invariant under the interchange of $n$ and $m$. As can be clearly seen, the three dominant Fourier components are $\cos(x)\cos(y)$, $\cos(3x)\cos(5y)$, and $\cos(5x)\cos(3y)$. According to the dispersion relation (13), the two latter modes have the same linear frequency $\sqrt{35\sqrt{34}}\approx 7.075\omega$, thus yielding a resonance of the seventh temporal harmonic. The eighth temporal harmonic, dominated by the $\cos(2x)\cos(6y)$ and $\cos(6x)\cos(2y)$ modes, is also nearly resonant, as indicated by linear frequency $\sqrt{41\sqrt{40}}\approx 7.975\omega$.
@@
Figure 16: Surface profiles of Case III standing waves for $k=1,l=0$. (a)-(c) A non-resonant solution with $H=0.3$ at $t=0,T/4$, and $T/2$. (d) A resonant solution with $H=0.29$ at $t=0$.
Figure 16 shows typical Case III standing waves for $k=1,l=0$. The surface profiles of a non-resonant solution with $H=0.3,\omega=1.3979$ are shown in panels (a)-(c), corresponding to $t=0,T/4$, and $T/2$. Note that this solution is essentially the same as the Case II standing wave with $k=l=\sqrt{2}/2,H=0.3$ after a rotation of the $(x,y)$-plane by $\pi/4$. This is also reflected by the fact that the shift connecting $\eta(x,y,0)$ and $\eta(x,y,T/2)$ is along the diagonal direction, which becomes the $x$- or $y$-direction after the rotation. Panel (d) exhibits a resonant standing wave with $H=0.29,\omega=1.3994$. The resonance is primarily between the fundamental and ninth temporal harmonics, which is dominated the $\cos(2x)\cos(5y)+\cos(5x)\cos(2y)$ modes.
@@
Figure 20: A Case II standing wave with $k=l=0.5,H=0.73$ on branch $1$. (a)-(c) show the top views of the solution at $t=0,T/8$ and $3T/16$. A periodic cell is surrounded by the dashed lines. The corresponding wave profiles are shown in (d)-(e).
Figure 21: A Case II standing wave with $k=l=0.5,H=0.19$ on branch $2$. (a)-(c) show the top views of the solution at $t=0,T/8$ and $3T/16$. A periodic cell is surrounded by the dashed lines. The corresponding wave profiles are shown in (d)-(e).
@@
where $a$ measures the relative importance of the two resonant modes, thus influencing the wave morphology. Panel (b) shows that $a$ takes opposite signs on the two branches and tends to $\pm 0.5$ as $\epsilon\rightarrow 0$. Figure 20 displays a representative standing wave with $H=0.73(\epsilon=0.4361),\omega=1.02800$ on branch $1$. Panels (a)-(c) show the top views of the solution at $t=0,T/8$ and $3T/16$, exhibiting waffle-shaped patterns clearly. The corresponding wave profiles within a periodic cell (the region surrounded by the dashed lines) are shown in panels (d)-(f). The solution is calculated using $128\times 128\times 128$ grid points. Similarly, figure 21 shows the top views and wave profiles of a branch-$2$ standing wave with $H=0.135(\epsilon=0.7095),\omega=0.98008$ at $t=0,T/8$ and $3T/16$. Figure 22 (a) and (b) exhibit the cross-sections of the two standing waves over a quarter temporal period, resembling the classical Wilton ripples when the fundamental and second-order Fourier components are resonant.
@@
Figure 25: A standing wave with $k=1,l=1.5312,H=0.4$ on branch $1$. (a)-(c) show the top views at $t=0,3T/16$ and $T/4$. The corresponding profiles are shown in (d)-(e) within the periodic cell surrounded by the dashed lines.
Figure 26: A standing wave with $k=1,l=1.5312,H=0.135$ on branch $2$. (a)-(c) show the top views at $t=0,3T/16$ and $T/4$. The corresponding three-dimensional profiles within a periodic cell (regions surrounded by the dashed lines) are shown in (d)-(e).
@@
At larger wave amplitude, branch-$1$ solutions exhibit similar wave profiles owing to the nearly constant value of $a$. Figure 25 shows the solution with $H=0.4(\epsilon=0.3480),\omega=2.7182$ on branch $1$. Panels (a)-(c) correspond to the top views of the solution at $t=0,3T/16$ and $T/4$, which clearly show the alternative appearance of hexagonal and quasi-two-dimensional patterns. The three-dimensional views of the wave profiles within a periodic cell are plotted in panels (d)-(f). In contrast, figure 26 exhibits the top views and wave profiles of a branch-$2$ standing wave with $H=0.135(\epsilon=0.3402),\omega=2.7222$. Both solutions are calculated using $96\times 96\times 96$ points. Figure 27 (a) and (b) show the cross-sections on the planes $y=0$ and $x=0$ for the branch-$1$ solution over a quarter temporal period. The former resembles the non-resonant two-dimensional standing wave profiles, while the latter is similar to the resonant profiles shown in figure 22.
Figure 28: Four bifurcations for standing waves with $k=0.459181,l=0.145664$. (a) $\omega$ versus $H$. (b)-(e) Amplitude ratio versus $\epsilon$.
Figure 29: Standing waves with $k=0.459181,l=0.145664$ at $t=0,T/8,T/4$. (a)-(c) A branch-11 solution with $H=0.2$. (d)-(f) A branch-22 solution with $H=0.07$. (g)-(i) A branch-33 solution with $H=0.195$. (j)-(l) A branch-44 solution with $H=0.185$. For the convenience of visualisation, the figures are stretches in the $x$- and $z$-directions.
@@
It is expected that there are four branches of solutions, corresponding to the combinations of the signs of $a_{1}$ and $a_{2}$: (+,+),(+,),(,+)(+,+),(+,-),(-,+), and (,)(-,-). Figure 28 (a) shows the frequency curves of the four bifurcations, all stemming from the infinitesimal linear solutions at the frequency $\omega=0.770405$. Branches $1,2$ and $4$ exhibit frequency down-shift, while branch $3$ shows a frequency up-shift. On the other hand, only three branches ($2,3$ and $4$) exhibit resonant standing waves. This can be viewed in panels (b)-(e) where we plot the amplitude ratios versus $\epsilon$ along the four bifurcations. For branch $1$, it is clear that $a_{1}\propto-\epsilon^{2}$ and $a_{2}\propto-\epsilon^{3}$ for $\epsilon\ll 1$. Consequently, the small-amplitude standing waves on this branch are non-resonant. In contrast, the amplitude ratios take much greater values on the other three branches. When $\epsilon\rightarrow 0$, $a_{1}\rightarrow-0.75,a_{2}\rightarrow 0.6$ on branch $2$, $a_{1}\rightarrow 0.75,a_{2}\rightarrow 0.6$ on branch $3$, and $a_{1}\rightarrow 0.75,a_{2}\rightarrow-0.6$ on branch $4$. No solution is found for the case $a_{1}\rightarrow-0.75,a_{2}\rightarrow-0.6$. Based on our numerical experiments using various initial guesses, it is believed that there are only four solution branches growing from the bifurcation point, excluding the trivia case that a single periodic wave repeats itself multiple times within one computational domain. The profiles at $t=0,T/4$ and $T/2$ of a non-resonant solution and three flower-like resonant solutions are shown in figure 29, where the four rows from top to bottom correspond to the branches $1$ to $4$, respectively.

### v2

@@ 1 Introduction
Two-dimensional standing water waves have been extensively investigated through rigorous analysis, weakly nonlinear theory, fully nonlinear simulations, and laboratory experiments. Plotnikov & Toland [plotnikov2001nash] and Iooss et al. [iooss2005standing] proved the existence of small-amplitude gravity standing waves in finite and infinite depth, respectively. The Stokes-type expansions of gravity standing waves were derived by Rayleigh [Rayleigh1915deep] and Penney & Price [penney1952some] for deep water, by Tadjbaksh & Keller [tadjbakhsh1960standing] for finite-depth water, and later extended to include surface tension by Concus [concus1962standing] and Vanden-Broeck [vanden1984nonlinear]. In recent years, direct numerical computations on two-dimensional standing waves have received more attention in studying the branching phenomenon [okamura1985instability, mercer1994form, smith1999branching, wilkening2012overdetermined, shelton2023structure], the mechanisms of instability [mercer1992standing, bryant1994different, wilkening2019harmonic], and the long-standing conjecture on the limiting configuration of gravity standing waves [mercer1992standing, schultz1998highly, okamura1998enclosed, okamura2003standing, okamura2010almost, wilkening2011breakdown, wilkening2012overdetermined]. Three principal numerical approaches have been established: (i) spatio-temporal collocation methods based on the Fourier and other spectral expansions [vanden1981numerical, tsai1987numerical, tsai1994numerical, bryant1994different, bryant1995water, okamura2003standing], (ii) semi-analytic series-expansion methods [schwartz1981semi, rottman1982steep, abassi2025semi], and (iii) time-dependent methods (usually via the boundary-integral formulation) combined with Newton or nonlinear least-squares solvers [mercer1992standing, schultz1998highly, wilkening2011breakdown, wilkening2012overdetermined, shelton2023structure]. The last approach is most commonly used owing to its relatively low computational cost, although the time integration may suffer from numerical instabilities, aliasing errors, and filtering errors. In contrast, the spatio-temporal collocation methods avoid these issues and guarantee exact temporal periodicity owing to the Fourier expansion in time. When the system is stiff or chaotic, this approach is particularly attractive, despite the heavy computational cost in solving large-scale algebraic systems. Such an issue had limited the application of spatio-temporal collocation methods in early years, but is largely alleviated over the past decades with a rapid growth of computational power.
@@ 2 Formulations of deep-water gravity-capillary waves
where $\mathcal{N}_{i}$ represents the DNO-involved $i$-th order nonlinear term whose expression can be found in Appendix 1. Wang & Milewski [wang2012dynamics] and Wang et al. [wang2014computation] employed a cubic ($n=3$) and a quintic ($n=5$) truncations (denoted by cubic model and quintic model hereafter) to study gravity-capillary and flexural-gravity solitary waves, respectively. They showed that these models are quite accurate compared with the original water-wave formulation, even when solutions have large amplitudes. In the following sections, we investigate three-dimensional gravity-capillary standing waves in rectangular or square basins by searching for triply periodic (in the $x,y$ and $t$ directions) solutions of the cubic and quintic models.
@@ 3.2 Spatio-temporal symmetries
    $\eta(x,y,T/4-t^{\prime})$ $=\eta(x\pm L_{1},y,T/4+t^{\prime})=\eta(x,y\pm L_{2},T/4+t^{\prime}),$ (22)
    $\varphi(x,y,T/4-t^{\prime})$ $=-\varphi(x\pm L_{1},y,T/4+t^{\prime})=-\varphi(x,y\pm L_{2},T/4+t^{\prime}).$ (23)
@@
    $\eta(x,y,T/4-t^{\prime})$ $=\eta(x\pm L_{1},y\pm L_{2},T/4+t^{\prime}),$ (24)
    $\varphi(x,y,T/4-t^{\prime})$ $=-\varphi(x\pm L_{1},y\pm L_{2},T/4+t^{\prime}).$ (25)
@@
    $\eta(-x,y,T/4)$ $=\eta(x\pm L_{1},y,T/4)=\eta(x,y\pm L_{2},T/4)=\eta(x,-y,T/4),$ (36)
    $\varphi(-x,y,T/4)$ $=-\varphi(x\pm L_{1},y,T/4)=-\varphi(x,y\pm L_{2},T/4)=\varphi(x,-y,T/4).$ (37)
@@ 3.3 Numerical implementation
$x_{i}$ $=\frac{2(i-1)L_{1}}{I}-L_{1},\,\quad i=1,2,\cdots,I+1,$ (40)
$y_{j}$ $=\frac{2(j-1)L_{2}}{J}-L_{2},\quad j=1,2,\cdots,J+1,$ (41)
$t_{n}$ $=\frac{(n-1)T}{N},\qquad\qquad n=1,2,\cdots,N+1,$ (42)
@@
All together, these give rise to (I/4+1)(J/2+1)N/2J/4=O(IJN/16)(I/4+1)(J/2+1)N/2-J/4=O(IJN/16) unknowns for $\eta$ and $\varphi$. For Case II and Case III, it is convenient to let $I=J$. The total number of unknowns for $\eta$ and $\varphi$ are reduced to $(I/4+1)^{2}N/2=O(I^{2}N/32)$ for Case II and $(I/4+1)(I/2+1)N/2=O(I^{2}N/16)$ for Case III. A schematic is drawn in figure 5 to show the reduced computational domain for all three cases. (10) and (11) give rise to the same number of equations after using the spatio-temporal symmetries. To close the system, we use the following wave-amplitude equation
@@ 4.2 Standing waves and bifurcations
We first show representative Case II standing waves with $k=l=1$. In contrast to the gravity standing waves featuring pyramidal-shaped crests, gravity-capillary standing waves exhibit rounded crests and troughs. Figure 8 displays a solution with $H=0.7,\omega=1.867$, calculated using $128\times 128\times 128$ grid points. Panels (a-c) correspond to the wave profiles at $t=0,T/4$ and $T/2$, respectively. At $t=T/4$, the surface develops multiple crests and troughs with the maximum wave amplitude decreasing to $0.16$. Figure 10 exhibits the energy distribution among the kinetic, gravitational, and surface-tension parts over one temporal period. The kinetic energy vanishes every half period, while the potential energy reaches its minimum, $0.497$, at $t=T/4,3T/4,\cdots$. The total energy remains at a constant level of $7.661$ perfectly, with the maximum relative error being $7\times 10^{-8}$.
@@
Figure 11: (a-e) Surface profiles of a resonant Case II standing wave with $k=l=1,H=0.33$ at $t=0,T/16,3T/16,7T/32$ and $T/4$. (f) Amplitudes of the first $100$ Fourier coefficients for $\eta(x,y,0)$.
A common feature for two-dimensional standing waves is the existence of various nonlinear resonant solutions [smith1999branching, wilkening2011breakdown, wilkening2012overdetermined]. We observe a similar phenomenon in three dimensions. Figure 11 shows a representative Case II resonant standing wave with $k=l=1,H=0.33,\omega=2.0192$, calculated using $128\times 128\times 128$ grid points. In contrast to the smooth surface profiles shown in figure 8, the resonant standing wave exhibits complex surface wrinkles. A local hump and dip alternately emerge on top of the crest at a faster oscillating frequency relative to the fundamental Fourier mode, as shown in panels (a-c). The surface develops a volcano-shaped appearance with four saddles along the diagonals at $t=7T/32$, as shown in panel (d). At $t=T/4$, the wave amplitude drops to $0.0247$ approximately, and the surface profile shows a similar pattern to that in figure 8 (b) but has much flatter crests and four sharper spikes at $(\pm\pi/2,\pm\pi/2)$. Panel (f) shows the Fourier amplitudes associated with $\cos(nx)\cos(my)$ mode for $\eta(x,y,0)$. Note that all Fourier modes with odd-valued sum $n+m$ vanish, in consequence of (22) and (23). Owing to the eightfold symmetry, the spectrum is invariant under the interchange of $n$ and $m$. As can be clearly seen, the three dominant Fourier components are $\cos(x)\cos(y)$, $\cos(3x)\cos(5y)$, and $\cos(5x)\cos(3y)$. According to the dispersion relation (13), the two latter modes have the same linear frequency $\sqrt{35\sqrt{34}}\approx 7.075\omega$, thus yielding a resonance of the seventh temporal harmonic. The eighth temporal harmonic, dominated by the $\cos(2x)\cos(6y)$ and $\cos(6x)\cos(2y)$ modes, is also nearly resonant, as indicated by linear frequency $\sqrt{41\sqrt{40}}\approx 7.975\omega$.
@@
Figure 15: Surface profiles of Case III standing waves for $k=1,l=0$. (a-c) A non-resonant solution with $H=0.3$ at $t=0,T/4$, and $T/2$. (d) A resonant solution with $H=0.29$ at $t=0$.
Figure 15 shows typical Case III standing waves for $k=1,l=0$. The surface profiles of a non-resonant solution with $H=0.3,\omega=1.3979$ are shown in panels (a-c), corresponding to $t=0,T/4$, and $T/2$. Note that this solution is essentially the same as the Case II standing wave with $k=l=\sqrt{2}/2,H=0.3$ after a rotation of the $(x,y)$-plane by $\pi/4$. This is also reflected by the fact that the shift connecting $\eta(x,y,0)$ and $\eta(x,y,T/2)$ is along the diagonal direction, which becomes the $x$- or $y$-direction after the rotation. Panel (d) exhibits a resonant standing wave with $H=0.29,\omega=1.3994$. The resonance is primarily between the fundamental and ninth temporal harmonics, which is dominated the $\cos(2x)\cos(5y)+\cos(5x)\cos(2y)$ modes.
@@
Figure 19: A Case II standing wave with $k=l=0.5,H=0.73$ on branch $1$. (a-c) show the top views of the solution at $t=0,T/8$ and $3T/16$. A periodic cell is surrounded by the dashed lines. The corresponding wave profiles are shown in (d) and (e).
Figure 20: A Case II standing wave with $k=l=0.5,H=0.19$ on branch $2$. (a-c) show the top views of the solution at $t=0,T/8$ and $3T/16$. A periodic cell is surrounded by the dashed lines. The corresponding wave profiles are shown in (d) and (e).
@@
where $a$ measures the relative importance of the two resonant modes, thus influencing the wave morphology. Panel (b) shows that $a$ takes opposite signs on the two branches and tends to $\pm 0.5$ as $\epsilon\rightarrow 0$. Figure 19 displays a representative standing wave with $H=0.73(\epsilon=0.4361),\omega=1.02800$ on branch $1$. Panels (a-c) show the top views of the solution at $t=0,T/8$ and $3T/16$, exhibiting waffle-shaped patterns clearly. The corresponding wave profiles within a periodic cell (the region surrounded by the dashed lines) are shown in panels (d-f). The solution is calculated using $128\times 128\times 128$ grid points. Similarly, figure 20 shows the top views and wave profiles of a branch-$2$ standing wave with $H=0.135(\epsilon=0.7095),\omega=0.98008$ at $t=0,T/8$ and $3T/16$. Figure 21 (a) and (b) exhibit the cross-sections of the two standing waves over a quarter temporal period, resembling the classical Wilton ripples when the fundamental and second-order Fourier components are resonant.
@@
Figure 24: A standing wave with $k=1,l=1.5312,H=0.4$ on branch $1$. (a-c) show the top views at $t=0,3T/16$ and $T/4$. The corresponding profiles within the periodic cell surrounded by the dashed lines are shown in (d) and (e).
Figure 25: A standing wave with $k=1,l=1.5312,H=0.135$ on branch $2$. (a-c) show the top views at $t=0,3T/16$ and $T/4$. The corresponding three-dimensional profiles within a periodic cell (regions surrounded by the dashed lines) are shown in (d) and (e).
@@
At larger wave amplitude, branch-$1$ solutions exhibit similar wave profiles owing to the nearly constant value of $a$. Figure 24 shows the solution with $H=0.4(\epsilon=0.3480),\omega=2.7182$ on branch $1$. Panels (a-c) correspond to the top views of the solution at $t=0,3T/16$ and $T/4$, which clearly show the alternative appearance of hexagonal and quasi-two-dimensional patterns. The three-dimensional views of the wave profiles within a periodic cell are plotted in panels (d-f). In contrast, figure 25 exhibits the top views and wave profiles of a branch-$2$ standing wave with $H=0.135(\epsilon=0.3402),\omega=2.7222$. Both solutions are calculated using $96\times 96\times 96$ points. Figure 26 (a) and (b) show the cross-sections on the planes $y=0$ and $x=0$ for the branch-$1$ solution over a quarter temporal period. The former resembles the non-resonant two-dimensional standing wave profiles, while the latter is similar to the resonant profiles shown in figure 21.
Figure 27: Four bifurcations for standing waves with $k=0.459181,l=0.145664$. (a) $\omega$ versus $H$. (b-e) Amplitude ratio versus $\epsilon$.
Figure 28: Standing waves with $k=0.459181,l=0.145664$ at $t=0,T/8,T/4$. (a-c) A branch-11 solution with $H=0.2$. (d-f) A branch-22 solution with $H=0.07$. (g-i) A branch-33 solution with $H=0.195$. (j-l) A branch-44 solution with $H=0.185$. For the convenience of visualisation, the figures are stretches in the $x$- and $z$-directions.
@@
It is expected that there are four branches of solutions, corresponding to the combinations of the signs of $a_{1}$ and $a_{2}$: (+,+),(+,),(,+)(+,+),(+,-),(-,+), and (,)(-,-). Figure 27 (a) shows the frequency curves of the four bifurcations, all stemming from the infinitesimal linear solutions at the frequency $\omega=0.770405$. Branches $1,2$ and $4$ exhibit frequency down-shift, while branch 33 shows a frequency up-shift. On the other hand, only three branches ($2,3$ and $4$) exhibit resonant standing waves. This can be viewed in panels (b-e) where we plot the amplitude ratios versus $\epsilon$ along the four bifurcations. For branch $1$, it is clear that $a_{1}\propto-\epsilon^{2}$ and $a_{2}\propto-\epsilon^{3}$ for $\epsilon\ll 1$. Consequently, the small-amplitude standing waves on this branch are non-resonant. In contrast, the amplitude ratios take much greater values on the other three branches. When $\epsilon\rightarrow 0$, $a_{1}\rightarrow-0.75,a_{2}\rightarrow 0.6$ on branch $2$, $a_{1}\rightarrow 0.75,a_{2}\rightarrow 0.6$ on branch $3$, and $a_{1}\rightarrow 0.75,a_{2}\rightarrow-0.6$ on branch $4$. No solution is found for the case $a_{1}\rightarrow-0.75,a_{2}\rightarrow-0.6$. Based on our numerical experiments using various initial guesses, it is believed that there are only four solution branches growing from the bifurcation point, excluding the trivia case that a single periodic wave repeats itself multiple times within one computational domain. The profiles at $t=0,T/4$ and $T/2$ of a non-resonant solution and three flower-like resonant solutions are shown in figure 28, where the four rows from top to bottom correspond to the branches $1$ to $4$, respectively.